OpenStudy (anonymous):

NEED HELP THNX :D Rewrite the following in the form f(x)=mx+b 1)1/3x+2y=-6 2)2x-4=3y+6 3)1/3y=2x+2/3

OpenStudy (jack1):

ok, so rearrange in terms of y = mx +c f(x) is sometime re-labelled as y c is a constant so Q1 1/3x+2y = -6 2y = -6 - 1/3x (subtract 1/3x from both sides) y = (-6 - 1/3x)/2 (divide both sides by 2) now u have the form they want... so answer to 1 y = -1/6x - 3 k now u try on number 2

OpenStudy (anonymous):

ok .. i think i got it .. 2x-4=3y+6  is it correct ?